The demand for company X's product is given by Qx = 2 - 3Px + 4Py

Suppose good X sells for $2.00 per unit and good Y sells for $4 per unit.

a. Calculate the cross-price elasticity of demand between goods X and Y at the given prices.

b. Are goods X and Y substitutes or complements?

c. What is the own price elasticity of demand at these prices?

Please show work
Answer #1

a) Ans: The cross-price elasticity of demand between goods X and Y at the given prices is 1.33.

Explanation:

Qx = 2 - 3Px + 4Py

= 2 - 3(2) + 4(4) = 2 - 6 + 16 = 12

Cross-price elasticity of demand = ∆Qx /∆Py * Py / Qx

= 4* (4/12) = 4 * 0.3333 = 1.33

Where , ∆Qx /∆Py = 4, which is the price coefficient of good Y.

b) Ans: Goods X and Y are substitutes.

Explanation:

When cross elasticity of demand is positive , then the two goods are substitutes.

When cross elasticity of demand is negative , then the two goods are complements.

c) Ans: The own price elasticity of demand at these prices is -0.5.

Explanation:

Elasticity of demand = ∆Qx /∆Px * Px / Qx

= -3 * ( 2 /12 )

= -3 * 0.1667

= -0.5

Where , ∆Qx /∆Px = -3 , which is the price coefficient of good X.